# Scheduling problem of aircrafts on a same runway and dual runways

### v5

@@ VII-B Algorithm 1
For the landing/takeoff scheduling problem, from Theorems 2, 5 and 9, the value of the objective function $F(\cdot)$ for the landing/takeoff scheduling problem is heavily related to the separation times between breakpoint aircraft and their trailing aircraft. To obtain or record all possible optimal sequences based on the bases of all the breakpoint-drift equivalent transformation sets, emphasis should be laid on the breakpoint aircraft. For the scheduling problem of landing and takeoff aircraft on a same runway, the number of breakpoint aircraft is usually small but the number of the takeoff-landing/landing-takeoff transitions might be large, and emphasis should be laid on the takeoff-landing/landing-takeoff transitions.
The goal of this step is actually to find all possible combinations of three consecutive aircraft of different classes in all sequences of the set $\Lambda_{i-1}$ to obtain the smallest value of the objective functions $F(\cdot)$, where each combination corresponds to a time increment when the aircraft $Tcf_{1i}$ is inserted. Starting from this goal, we can reduce some computation amount by giving up some unnecessary operations. For example, since we are concerned about only the time increment when the aircraft $Tcf_{1i}$ is inserted into $\phi_{(i-1)k}$, we need not to repeatedly consider the insertion of aircraft $Tcf_{1i}$ into the same class combination of aircraft and the same separation times, and only focus on the different class combination of aircraft or different separation times, which might significantly reduce the computation amount.
@@
$(i-2)$. Construct a sequence set $F_{inc}=\{\langle h_{1},h_{2},h_{3}\rangle\mid h_{1},h_{2},h_{3}\in\{1,2,\cdots,i-1\}\}$ composed of all the sequences of length $3$ such that
$\begin{array}[]{lll}Y_{pk}+Y_{k(1i)}+Y_{(1i)j}-Y_{pj}<f^{i}_{inc}-F(\phi_{i-1})\end{array}$ (20)
for any $\langle p,k,j\rangle\in F_{inc}$. Note here that since the minimum separation times are only related to the classes and the operation tasks of the aircraft, we can classify the aircraft to form the set $F_{inc}$ by calculating the inequality (20) according to the classes and the operation tasks of the aircraft.
@@
$(i-3).$ According to the elements of the set $F_{inc}$, adjust the aircraft orders in the sequence $\phi_{(i-1)k}$ and insert the aircraft $Tcf_{1i}$ to generate a new sequence $\bar{\phi}_{ik}$ under the condition $F(\bar{\phi}_{ik})<f^{i}_{inc}$ such that $\bar{\phi}_{ik}$ contains the subsequence $\langle Tcf_{p_{1}},Tcf_{k_{1}},Tcf_{1i},Tcf_{j_{1}}\rangle$ and $\langle p_{1},{k_{1}},{j_{1}}\rangle\in F_{inc}$. Minimize $F(\bar{\phi}_{ik})$ with the subsequence $\langle Tcf_{p_{1}},Tcf_{k_{1}},Tcf_{1i},Tcf_{j_{1}}\rangle$ unchanged based on the obtained theoretical results. Compare the values of all possible $F(\bar{\phi}_{ik})$ so as to find the optimal sequence of aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1i}$.
@@
To illustrate this step in a more specific way, consider a group of landing aircraft with $T_{0}=60$ and $\delta=8$ under the minimum separation time standards at Heathrow Airport. From the previous analysis in (i-2), the inequality (20) is equivalent to
$Y_{k(1i)}+Y_{(1i)j}-Y_{kj}<f^{i}_{inc}-F(\phi_{i-1}).$ (21)
@@
If $f^{i}_{inc}-F(\phi_{i-1})=T_{0}$, by simple calculations, the inequality (21) holds when $(cl_{k},cl_{1i},cl_{j})=(2,5,4),(3,6,5),(4,6,5)$. Clearly, when $(cl_{k},cl_{1i},cl_{j})=(2,5,4),(3,6,5),(4,6,5)$, each aircraft $Tcf_{k}$ is a breakpoint aircraft and the aircraft $Tcf_{1i}$ is needed to be inserted to be between aircraft $Tcf_{k}$ and $Tcf_{j}$. To generate a subsequence $\langle Tcf_{k},Tcf_{1i},Tcf_{j}\rangle$, we can focus on the breakpoint aircraft and their trailing aircraft, and based on the theoretical results in previous sections, by adjusting the aircraft orders properly, the optimal sequence can be found under the condition that $F(\bar{\phi}_{ik})<f^{i}_{inc}$ in an easy way. If $f^{i}_{inc}-F(\phi_{i-1})=T_{0}+\delta$, the generation of a subsequence $\langle Tcf_{k_{1}},Tcf_{1i},Tcf_{j_{1}}\rangle$ with $\langle{k_{1}},{j_{1}}\rangle\in F_{inc}$ would result in a new breakpoint aircraft or changes of the separation times between existing breakpoint aircraft and their trailing aircraft which might increase the value of the objective function $F(\cdot)$. Note for the separation times between breakpoint aircraft and their trailing aircraft that $T_{45}=T_{0}+\delta$ and $T_{ij}\geq 1.5T_{0}$ for all other $1\leq i<j\leq 6$. We can generate the whole sequence based on the subsequence $\langle Tcf_{k_{1}},Tcf_{1i},Tcf_{j_{1}}\rangle$ according to time increments of the new breakpoint aircraft or changes of the separation times between breakpoint aircraft and their trailing aircraft under the condition that $F(\bar{\phi}_{ik})<f^{i}_{inc}$, and the optimal sequence can also be found in an easy way. Besides these two cases, the case of $f^{i}_{inc}-F(\phi_{i-1})>T_{0}+\delta$ as well as the takeoff sequences and mixed landing and takeoff sequences on a same runway can be discussed in a similar easy way.
@@
From the above illustrations, we can first move the aircraft $Tcf_{p_{1}},Tcf_{k_{1}},Tcf_{j_{1}}$ to proper places by equivalent transformations including breakpoint-drift equivalent transformations and then use the extraction operation and insertion operation to generate the sequence $\bar{\phi}_{ik}$ with a focus on the breakpoint aircraft, their trailing aircraft and the landing-takeoff/takeoff-landing transitions as well as the properties of class-monotonically-decreasing sequences on the basis of the optimal sequence.
@@
Since $Y_{p_{1}k_{1}}+Y_{k_{1}(1i)}+Y_{(1i)j_{1}}-Y_{p_{1}j_{1}}<f^{i}_{inc}-F(\phi_{i-1})$, there is no utilizable subsequence $\langle Tcf_{p_{1}},Tcf_{k_{1}},Tcf_{j_{1}}\rangle$ in each $\phi_{(i-1)k}$ for $Tcf_{1i}$ to be inserted between aircraft $Tcf_{k_{1}}$ and $Tcf_{j_{1}}$. So, the generation of a subsequence $\langle Tcf_{p_{1}},Tcf_{k_{1}},Tcf_{j_{1}}\rangle$ would increase the value of the objective function $F(\cdot)$ for aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1(i-1)}$. Moreover, the optimal insertion of $Tcf_{1(i-1)}$ might usually generate a new breakpoint/landing-takeoff/takeoff-landing transition or change an existing separation time between a breakpoint aircraft and its trailing aircraft with some necessary aircraft order adjustments. These properties and related discussions in previous sections can be used to reduce the computation amount of Algorithm 1.
@@ VII-C Algorithm 2
Step 3. (3.1) Match the takeoff aircraft with each landing sequence in $\Lambda_{a}$ to generate a new sequence, denoted by $\Phi_{c}=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n+m}\rangle$, and optimize $\Phi_{c}$ such that $F(\Phi^{c}_{LA},Sr_{LA}(\Phi_{c}))=F(\Phi_{a})$ and $n_{z}$ is minimized, where $\Phi^{c}_{LA}$ and $Sr_{LA}(\Phi_{c})$ denote the landing sequence and the corresponding landing time vector in $\Phi_{c}$, and $n_{z}$ denotes the number of the aircraft of $\Phi_{c}$ which take off in $(t_{0}+F(\Phi_{a}),+\infty)$.
@@
(3.2) Insert the last $n_{z}$ aircraft of $\Phi_{c}$ into the subsequence of $\Phi_{c}$, $\Phi^{sub}_{c}=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n+m-n_{z}}\rangle$, to generate a new sequence $\Phi_{c3}$ and minimize $F(\Phi_{c3})$ without changing the orders of the landing aircraft based on the sets defined Step 2 by combing the length, the initial redundant time and the landing/takeoff time increment of each block subsequence. Note here that the sequence $\Phi_{c3}$ can contain D-block subsequences.
@@
Step 4. (4.1) The main idea of this step is to generate a solution $(\Phi_{c4},Sr(\Phi_{c4}))$ such that $\Phi^{c4}_{LA}\in\Lambda_{a}$, and search for the optimal solutions for the optimization problem (6) under the condition that $F(\Phi_{a})\leq F(\Phi^{c4}_{LA})\leq F(\Phi_{c3})$. Note that the set $\Lambda_{a}$ used here does not contain the information of the landing/takeoff times.
Since different D-block subsequences might have different landing time increments, we can search for the optimal solutions for the optimization problem (6) according to the sum of the landing time increments of D-block subsequences of the sequence $\Phi_{c4}$.
@@
When D-block subsequences of $\Phi_{c4}$ are considered, some necessary aircraft order adjustments might be needed to be imposed on the sequence $\Phi_{c4}$ to fully consider all the possibilities of the block subsequences to reduce the value of the objective function $F(\cdot)$. For example, suppose that $\phi_{LA}^{0}=\langle Tcf_{1},Tcf_{2},Tcf_{3},Tcf_{4},Tcf_{5}\rangle$ is a landing subsequence of a sequence $\phi_{0}$, and $\phi_{0}$ is composed of a D-Block subsequence $DB_{1}$ and a T-Block subsequence $TB_{1}$ with $Tcf_{1},Tcf_{2}\in DB_{1}$ and $Tcf_{2},Tcf_{3},Tcf_{4},Tcf_{5}\in TB_{1}$. According to the definition of D-block subsequence, $S_{12}>Y_{12}$. If $Y_{15}>Y_{12}$ and the objective function value of the landing sequence can be decreased by exchanging the orders of $Tcf_{2}$ and $Tcf_{5}$, the D-block subsequences of the sequence $\phi=\langle Tcf_{1},Tcf_{5},Tcf_{3},Tcf_{4},Tcf_{2}\rangle$ are needed to be considered.
@@
(4.2) Repeat Step (4.1) until all possible cases have been considered and the optimal sequence is found.
@@
The main idea of Algorithm 2 is to decompose the aircraft sequence into several block subsequences and fully explore combinations of the block subsequences along an optimal landing/takeoff subsequence to consider the optimization problem (6). Note that the number of all possible block subsequences of different class combinations is not large according to the minimum separation time standards at Heathrow Airport and for the RECAT-EU system, when resident-point aircraft are not taken into account. The use of block subsequences might significantly reduce the computation amount of the algorithm.

### v7

@@ VII-B Algorithm 1
For the landing/takeoff scheduling problem, from Theorems 2, 5 and 10, the value of the objective function $F(\cdot)$ is heavily related to the separation times between breakpoint aircraft and their trailing aircraft. To obtain or record all possible optimal sequences based on the bases of all the breakpoint-drift equivalent transformation sets, emphasis should be laid on the breakpoint aircraft. For the scheduling problem of landing and takeoff aircraft on a same runway, the number of breakpoint aircraft is usually small but the number of the takeoff-landing/landing-takeoff transitions might be large, and emphasis should be laid on the takeoff-landing/landing-takeoff transitions.
The goal of this step is actually to find all possible combinations of four consecutive aircraft of different classes in all sequences of the set $\Lambda_{i-1}$ to obtain the smallest value of the objective functions $F(\cdot)$, where each combination corresponds to a time increment when the aircraft $Tcf_{1i}$ is inserted. Starting from this goal, we can reduce some computation amount by giving up some unnecessary operations. For example, since we are concerned about only the time increment when the aircraft $Tcf_{1i}$ is inserted into $\phi_{(i-1)k}$, we need not to repeatedly consider the insertion of aircraft $Tcf_{1i}$ into the same class combination of aircraft and the same separation times, and focus on only the different class combination of aircraft or different separation times, which might significantly reduce the computation amount.
@@
$(i-2)$. Construct a sequence set $F_{inc}=\{\langle h_{1},h_{2},h_{3},h_{4}\rangle\mid h_{1},h_{2},h_{3},h_{4}\in\{1,2,\cdots,i-1\}\}$ composed of all the sequences of length $4$ such that
$\begin{array}[]{lll}\omega_{min}(\langle Tcf_{h_{1}},Tcf_{h_{2}},Tcf_{h_{3}},Tcf_{h_{4}}\rangle,Tcf_{1i})\\ <f^{i}_{inc}-F(\phi_{i-1})\end{array}$ (20)
for any $\langle h_{1},h_{2},h_{3},h_{4}\rangle\in F_{inc}$. Note here that since the minimum separation times are related to only the classes and the operation tasks of the aircraft, we can classify the aircraft to form the set $F_{inc}$ by calculating the inequality (20) according to the classes and the operation tasks of the aircraft.
@@
$(i-3).$ According to the elements of the set $F_{inc}$, adjust the aircraft orders in the sequence $\phi_{(i-1)k}$ and insert the aircraft $Tcf_{1i}$ to generate a new sequence $\bar{\phi}_{ik}$ under the condition $F(\bar{\phi}_{ik})<f^{i}_{inc}$ such that $\bar{\phi}_{ik}$ contains the subsequence $\langle Tcf_{h_{1}},Tcf_{h_{2}},Tcf_{1i},Tcf_{h_{3}},Tcf_{h_{4}}\rangle$ and $\langle h_{1},h_{2},h_{3},h_{4}\rangle\in F_{inc}$. Minimize $F(\bar{\phi}_{ik})$ based on the obtained theoretical results. Compare the values of all possible $F(\bar{\phi}_{ik})$ so as to find the optimal sequence of aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1i}$.
@@
To illustrate this step in a more specific way, consider a group of landing aircraft with $T_{0}=60$ and $\delta=8$ under the minimum separation time standards at Heathrow Airport. From the previous analysis in (i-2), the inequality (20) is equivalent to the form
$Y_{p(1i)}+Y_{(1i)j}-Y_{pj}<f^{i}_{inc}-F(\phi_{i-1}).$ (21)
@@
If $f^{i}_{inc}-F(\phi_{i-1})=T_{0}$, by simple calculations, the inequality (21) holds when $(cl_{p},cl_{1i},cl_{j})=(2,5,4),(3,6,5),(4,6,5)$. Clearly, when $(cl_{p},cl_{1i},cl_{j})=(2,5,4),(3,6,5),(4,6,5)$, each aircraft $Tcf_{p}$ is a breakpoint aircraft and the aircraft $Tcf_{1i}$ is needed to be inserted to be between aircraft $Tcf_{p}$ and $Tcf_{j}$. To generate a subsequence $\langle Tcf_{p},Tcf_{1i},Tcf_{j}\rangle$, we can focus on the breakpoint aircraft and their trailing aircraft, and based on the theoretical results in previous sections, by adjusting the aircraft orders properly, the optimal sequence can be found under the condition that $F(\bar{\phi}_{ik})<f^{i}_{inc}$ in an easy way. If $f^{i}_{inc}-F(\phi_{i-1})=T_{0}+\delta$, the generation of a subsequence $\langle Tcf_{p_{1}},Tcf_{1i},Tcf_{j_{1}}\rangle$ with $\langle{p_{1}},{j_{1}}\rangle\in F_{inc}$ would result in a new breakpoint aircraft or changes of the separation times between existing breakpoint aircraft and their trailing aircraft which might increase the value of the objective function $F(\cdot)$. Note for the separation times between breakpoint aircraft and their trailing aircraft that $T_{45}=T_{0}+\delta$ and $T_{ij}\geq 1.5T_{0}$ for all other $1\leq i<j\leq 6$. We can generate the whole sequence based on the subsequence $\langle Tcf_{p_{1}},Tcf_{1i},Tcf_{j_{1}}\rangle$ according to time increments of the new breakpoint aircraft or changes of the separation times between breakpoint aircraft and their trailing aircraft under the condition that $F(\bar{\phi}_{ik})<f^{i}_{inc}$, and the optimal sequence can also be found in an easy way. Besides these two cases, the case of $f^{i}_{inc}-F(\phi_{i-1})>T_{0}+\delta$ as well as the takeoff sequences and mixed landing and takeoff sequences on a same runway can be discussed in a similar easy way.
@@
From the above illustrations, we can first move the aircraft $Tcf_{h_{1}},Tcf_{h_{2}},Tcf_{h_{3}},Tcf_{h_{4}}$ to proper places by equivalent transformations including breakpoint-drift equivalent transformations and then use the extraction operation and insertion operation to generate the sequence $\bar{\phi}_{ik}$ with a focus on the breakpoint aircraft, their trailing aircraft and the landing-takeoff/takeoff-landing transitions as well as the properties of class-monotonically-decreasing sequences on the basis of the optimal sequence.
@@
Since each $\phi_{(i-1)k}$ is an optimal sequence for aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1(i-1)}$, if there is no utilizable subsequence $\langle Tcf_{h_{1}},Tcf_{h_{2}},Tcf_{h_{3}},Tcf_{h_{4}}\rangle$ formed by four consecutive aircraft of $Tcf_{h_{1}},Tcf_{h_{2}},Tcf_{h_{3}},Tcf_{h_{4}}$ in all $\phi_{(i-1)k}$, the generation of $\langle Tcf_{h_{1}},Tcf_{h_{2}},Tcf_{h_{3}},Tcf_{h_{4}}\rangle$ in $\phi_{(i-1)k}$ would increase the value of the objective function $F(\cdot)$, which is the reason why only the aircraft combinations satisfying (20) are taken into account. Moreover, the optimal insertion of $Tcf_{1(i-1)}$ might usually generate a new breakpoint/landing-takeoff/takeoff-landing transition or change an existing separation time between a breakpoint aircraft and its trailing aircraft with some necessary aircraft order adjustments. These properties and related discussions in previous sections can be used to reduce the computation amount of Algorithm 1.
Based on the above analysis, we can discuss the optimality of the sequence $\bar{\phi}_{ik}$ in the following cases.
Scenario 1. Landing/takeoff sequences. Let $Nb_{\bar{\phi}_{ik}}$ and $Nb_{\phi_{(i-1)k}}$ denotes the numbers of breakpoint aircraft in the sequences $\bar{\phi}_{ik}$ and $\phi_{(i-1)k}$.
Case 1.1. $Nb_{\bar{\phi}_{ik}}\leq Nb_{\phi_{(i-1)k}}$.
Case 1.2. $Nb_{\bar{\phi}_{ik}}>Nb_{\phi_{(i-1)k}}$.
Scenario 2. Mixed landing and takeoff sequences on a same runway. Let $Nt_{\bar{\phi}_{ik}}$ and $Nt_{\phi_{(i-1)k}}$ denotes the numbers of landing-takeoff and takeoff-landing transitions in the sequences $\bar{\phi}_{ik}$. and $\phi_{(i-1)k}$.
Case 2.1. $Nt_{\bar{\phi}_{ik}}\leq Nt_{\phi_{(i-1)k}}$.
Subcase 2.1.1. $Nb_{\bar{\phi}_{ik}}\leq Nb_{\phi_{(i-1)k}}$.
Subcase 2.1.2. $Nb_{\bar{\phi}_{ik}}>Nb_{\phi_{(i-1)k}}$.
Case 2.2. $Nt_{\bar{\phi}_{ik}}>Nt_{\phi_{(i-1)k}}$.
Subcase 2.2.1. $Nb_{\bar{\phi}_{ik}}\leq Nb_{\phi_{(i-1)k}}$.
Subcase 2.2.2. $Nb_{\bar{\phi}_{ik}}>Nb_{\phi_{(i-1)k}}$.
When the number of breakpoint aircraft in $\bar{\phi}_{ik}$ is 2 larger than that in $\phi_{(i-1)k}$, $\bar{\phi}_{ik}$ is usually not the optimal sequence when the aircraft subsequence $\langle Tcf_{i},Tcf_{j}\rangle$ with $cl_{i}=\rho_{2}-1$ and $cl_{j}=\rho_{2}$ is not involved.
Note that $\phi_{(i-1)k}$ might be composed of multiple class-monotonically-decreasing subsequences, denoted by $\phi^{1},\phi^{2},\cdots,\phi^{s}$. We can consider the insertion of $Tcf_{1i}$ on each subsequence $\phi^{k}$ with some corresponding aircraft order adjustments so as to reduce the complexity of the analysis of the optimality of the sequence.
@@ VII-C Algorithm 2
Step 3. (3.1) Match the takeoff aircraft with each landing sequence in $\Lambda_{a}$ to generate a new sequence, denoted by $\Phi_{c}=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n+m}\rangle$, and optimize $\Phi_{c}$ such that $F(\Phi^{c}_{LA},Sr_{LA}(\Phi_{c}))=F(\Phi_{a})$, where $\Phi^{c}_{LA}$ and $Sr_{LA}(\Phi_{c})$ denote the landing sequence and the corresponding landing time vector in $\Phi_{c}$.
@@
(3.2) Let $n_{z}$ denote the number of the aircraft of $\Phi_{c}$ which take off in $(t_{0}+F(\Phi_{a}),+\infty)$. Insert the last $n_{z}$ aircraft of $\Phi_{c}$ into the subsequence of $\Phi_{c}$, $\Phi^{sub}_{c}=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n+m-n_{z}}\rangle$, to generate a new sequence $\Phi_{c3}$ and minimize $F(\Phi_{c3})$ without changing the orders of the landing aircraft based on the sets defined Step 2 by combing the length, the initial redundant time and the landing/takeoff time increment of each block subsequence. Note here that the sequence $\Phi_{c3}$ can contain D-block subsequences.
@@
Step 4. The main idea of this step is to implement a full space search for an optimal solution $(\Phi_{c4},Sr(\Phi_{c4}))$ for the optimization problem (6) under the condition that $F(\Phi_{a})\leq F(\Phi_{c4})\leq F(\Phi_{c3})$ starting from the initial condition that $\Phi^{c4}_{LA}\in\Lambda_{a}$. Note that the set $\Lambda_{a}$ used here does not contain the information of the landing/takeoff times.
Since different D-block subsequences might have different landing time increments, we can search for the optimal solutions for the optimization problem (6) by increasing the sum of the landing time increments of D-block subsequences of the sequence $\Phi_{c4}$. Based on this idea, since $F(\Phi_{a})\leq F(\Phi^{c4}_{LA})\leq F(\Phi_{c4})$, the condition $F(\Phi_{a})\leq F(\Phi_{c4})\leq F(\Phi_{c3})$ can be relaxed to be $F(\Phi_{a})\leq F(\Phi^{c4}_{LA})\leq F(\Phi_{c3})$.
@@
When D-block subsequences of $\Phi_{c4}$ are considered, some necessary aircraft order adjustments might be needed to be imposed on the subsequence $\Phi^{c4}_{LA}$ to fully consider all the possibilities of the block subsequences to reduce the value of the objective function $F(\cdot)$. For example, suppose that $\phi_{LA}^{0}=\langle Tcf_{1},Tcf_{2},Tcf_{3},Tcf_{4},Tcf_{5}\rangle$ is a landing subsequence of a sequence $\phi_{0}$, and $\phi_{0}$ is composed of a D-Block subsequence $DB_{1}$ and a T-Block subsequence $TB_{1}$ with $Tcf_{1},Tcf_{2}\in DB_{1}$ and $Tcf_{2},Tcf_{3},Tcf_{4},Tcf_{5}\in TB_{1}$. According to the definition of D-block subsequence, $S_{12}>Y_{12}$. If $Y_{15}>Y_{12}$ and the objective function value of the landing sequence can be decreased by exchanging the orders of $Tcf_{2}$ and $Tcf_{5}$, the D-block subsequences of the sequence $\phi=\langle Tcf_{1},Tcf_{5},Tcf_{3},Tcf_{4},Tcf_{2}\rangle$ are needed to be considered.
@@
The main idea of Algorithm 2 is to decompose the aircraft sequence into several block subsequences and fully explore combinations of the block subsequences starting from an optimal landing/takeoff subsequence to consider the optimization problem (6). Note that the number of all possible block subsequences of different class combinations is not large according to the minimum separation time standards at Heathrow Airport and for the RECAT-EU system, when resident-point aircraft are not taken into account. The use of block subsequences might significantly reduce the computation amount of the algorithm.